# Bounded geometry for PCF-special subvarieties

### v2

@@ 3.2. Measures on $\mathcal{X}_{V}^{\ell+1}$
As $\alpha_{n_{0}}(0)\in\mathbb{D}_{\varepsilon/2}$, the inequality (4.9) gives $|\nabla\alpha_{n_{0}}(x)|<M_{2}\leq M_{0}$, so that $\alpha_{n_{0}}(x)$ is in $\mathbb{D}_{\varepsilon}$ for each $x\in\mathbb{D}^{r}_{\delta}$. By induction, we see that if $\alpha_{n_{0}+i-1}(x)$ is in $\mathbb{D}_{\varepsilon}$ with $|\nabla\alpha_{n_{0}+i-1}|\leq M_{i+1}$, then $\alpha_{n_{0}+i}(x)$ is in $\mathbb{D}_{\varepsilon}$, because $\alpha_{n_{0}+i}(0)$ is in $\mathbb{D}_{\varepsilon/2}$ and $\delta<\varepsilon/(2M_{0})$. Since $M_{i}\to 0$ as $i$ tends to infinity, we see that $\alpha_{n}$ converges to 00 uniformly on $\mathbb{D}^{r}_{\delta}$. ∎
We now return to our proof. Abusing the notation slightly, we write $\{f_{s}\}_{s\in X_{\lambda_{1}}}$ to denote the restriction of $f_{\ell}$ on $X_{\lambda_{1}}$, $c_{r+1}$ for the restriction of $c_{r+1}$ to $X_{\lambda_{1}}$, and $T_{c_{r+1}}$ its associated bifurcation current on $X_{\lambda_{1}}$. We know from Lemma 4.1 that, zooming into $s_{\ell}$ in $X_{\lambda_{1}}$ at any exponential rate, the graphs of iterates of $c_{r+1}$ are converging to the graph of a constant function.
Let $p_{0}\in\mathbb{P}^{1}$ be the parabolic point of $f_{s_{\ell}}$; replacing $f_{s_{\ell}}$ with an iterate, we have $f_{s_{\ell}}^{n}(c_{r+1}(s_{\ell}))\to p_{0}$ as $n\to\infty$. After a change of coordinates if necessary, we may assume that $c_{r+1}(s_{\ell})$ and $p_{0}$ are not $\infty$. Recalling the definition of the bifurcation current associated to $c_{r+1}$, we can work in homogeneous coordinates on $\mathbb{P}^{1}$ over a neighborhood of $s_{\ell}$ in $X_{\lambda_{1}}$. We define
@@
Following the notation of [Ga2, Proposition 3.2], we now let $r_{n}:\mathbb{D}^{r}\to X_{\lambda_{1}}$ be the sequence of maps that define the zooms near the parameter $s_{\ell}$ for the measure $\mu^{r+1}_{\lambda_{1}}$, as described before Lemma 4.1. It follows from (4.10) that the functions $U_{n}\circ r_{n}$ also converge uniformly to 0. We conclude that $d^{n}\,(r_{n})^{*}T_{c_{r+1}}\to 0$ weakly, and therefore
@@
We have thus reached our desired contradiction, and this completes the proof of Theorem 1.8.
@@
The proof would simplify if we knew that the $(r+1)$-th critical point $c_{r+1}$ were preperiodic to a parabolic cycle, instead of only in the basin of that cycle. We could employ ideas that appeared in [GNY, Proposition 5.2] (written for measures in the dynamical plane) or [FG3, Proposition 5.12] (for $1$-parameter families of polynomials and marked points). But the critical point $c_{r+1}$ is sometimes the only “free” critical point and must therefore have an infinite orbit in the presence of a parabolic cycle.
@@
We then repeat the argument with the families of curves over the positive-dimensional components in the complement of $V\setminus U$; we work with irreducible components so that the generic fibers of these families will be (geometrically) irreducible. Note that each family we treat is defined over $\overline{\mathbb{Q}}$. We increase the bound $N$ and decrease $\epsilon$ as needed. We are left with a finite set $S$ in $V(\overline{\mathbb{Q}})$ containing all of the irreducible PCF-special curves with degree $\leq D$ as well as all curves in $V$ with infinitely many points of critical height $<\frac{\varepsilon}{\ell+1}\max\{1,h_{\mathrm{Ch}}(\lambda)\}$ (or $<\frac{\varepsilon}{\ell+1}$ over complex parameters). This proves Theorems 1.2 and 1.6 for the curves in $V$. If for any of the curves $C_{\lambda}$ in this leftover finite set $S\subset V(\overline{\mathbb{Q}})$, we have $N<|\Sigma_{\lambda,\varepsilon}|<\infty$, we just increase the bound $N$ as needed so that Theorem 1.5 and 1.2 hold for all curves in $V$. We repeat the argument for all irreducible components $V$ of $\mathrm{Ch}(\mathrm{M}_{d},1,D)$ with geometrically irreducible generic fiber and adjust the constants $\varepsilon,~{}N$ as needed. This proves Theorems 1.1, 1.2, 1.5 and 1.6 for curves.
Now fix $r>1$ and suppose we know the conclusion of Theorems 1.1, 1.2, 1.5, and 1.6 for subvarieties in $\mathrm{M}_{d}$ of dimension $<r$. Consider any family $\mathcal{X}\to V$ of $r$-dimensional varieties in $\mathrm{M}_{d}$ over an irreducible and quasiprojective $V\subset\mathrm{Ch}(\mathrm{M}_{d},r,D)$ defined over $\overline{\mathbb{Q}}$ and of dimension $\ell\geq 1$ for which the generic fiber of $\mathcal{X}\to V$ is geometrically irreducible.
@@
for $\lambda\in U(\mathbb{C})\setminus U(\overline{\mathbb{Q}})$, lie in a proper algebraic subvariety $Z_{(\lambda)}\subset X_{\lambda}$ which has $\leq N_{1}$ irreducible components, each with degree $\leq B_{1}$, for all $\lambda\in U(\mathbb{C})$. Note that $\Sigma^{r}_{\lambda,\epsilon}$ is not necessarily Zariski dense in $Z_{(\lambda)}$, so we apply the induction hypotheses (specifically, the constants from each of the statements of Theorems 1.1, 1.2, 1.5, 1.6 for subvarieties of dimension $<r$) to each irreducible component $Y_{(\lambda)}$ of $Z_{(\lambda)}$, for all $\lambda\in U(\mathbb{C})$. Thus, by reducing $\varepsilon$ if needed, we conclude that the Zariski-closure of $\Sigma^{r}_{\lambda,\varepsilon}\cap Y_{(\lambda)}$ has uniformly bounded number of components, each with uniformly bounded degree for all $\lambda\in U(\mathbb{C})$ Further, adjusting these uniform bounds if needed, using the induction hypothesis, we can arrange that they also bound the geometry of the Zariski closure of $\Sigma^{r}_{\lambda,\varepsilon^{\prime}}\cap Y_{(\lambda)}\subset\Sigma^{r}_{\lambda,\varepsilon}\cap Y_{(\lambda)}$ for all $0\leq\varepsilon^{\prime}<\varepsilon$ (and in particular of $Y_{(\lambda)}\cap\mathrm{PCF}$). Combined with the initial bound of $N_{1}$ on the number of components of $Z_{(\lambda)}$, repeating this procedure provides uniform bounds on the Zariski closures of $\Sigma^{r}_{\lambda,\varepsilon^{\prime}}$ for all $\varepsilon^{\prime}<\varepsilon$ and of $\overline{X_{\lambda}\cap\mathrm{PCF}}$ for all $\lambda\in U(\mathbb{C})$.

### v3

@@ 3.2. Measures on $\mathcal{X}_{V}^{\ell+1}$
As $\alpha_{n_{0}}(0)\in\mathbb{D}_{\varepsilon/2}$, the inequality (4.9) gives $|\nabla\alpha_{n_{0}}(x)|<M_{2}\leq M_{0}$, so that $\alpha_{n_{0}}(x)$ is in $\mathbb{D}_{\varepsilon}$ for each $x\in\mathbb{D}^{r}_{\delta}$. By induction, we see that if $\alpha_{n_{0}+i-1}(x)$ is in $\mathbb{D}_{\varepsilon}$ with $|\nabla\alpha_{n_{0}+i-1}|\leq M_{i+1}$, then $\alpha_{n_{0}+i}(x)$ is in $\mathbb{D}_{\varepsilon}$, because $\alpha_{n_{0}+i}(0)$ is in $\mathbb{D}_{\varepsilon/2}$ and $\delta<\varepsilon/(2M_{0})$. Since $M_{i}\to 0$ as $i$ tends to infinity, we see that $\alpha_{n}$ converges to 0 uniformly on $\mathbb{D}^{r}_{\delta}$. ∎
We now return to our proof. Abusing the notation slightly, we write $\{f_{s}\}_{s\in X_{\lambda_{1}}}$ to denote the restriction of the family of maps to $X_{\lambda_{1}}$, $c_{r+1}$ for the restriction of $c_{r+1}$ to $X_{\lambda_{1}}$, and $T_{c_{r+1}}$ its associated bifurcation current on $X_{\lambda_{1}}$. We know from Lemma 4.1 that, zooming into $s_{\ell}$ in $X_{\lambda_{1}}$ at any exponential rate, the graphs of iterates of $c_{r+1}$ are converging to the graph of a constant function.
Let $p_{0}\in\mathbb{P}^{1}$ be the parabolic periodic point of $f_{s_{\ell}}$; replacing $f_{s_{\ell}}$ with an iterate, we have $f_{s_{\ell}}^{n}(c_{r+1}(s_{\ell}))\to p_{0}$ as $n\to\infty$. After a change of coordinates if necessary, we may assume that $c_{r+1}(s_{\ell})$ and $p_{0}$ are not $\infty$. Recalling the definition of the bifurcation current associated to $c_{r+1}$, we can work in homogeneous coordinates on $\mathbb{P}^{1}$ over a neighborhood of $s_{\ell}$ in $X_{\lambda_{1}}$. We define
@@
Following the notation of [Ga2, Proposition 3.2], we now let $r_{n}:\mathbb{D}^{r}\to X_{\lambda_{1}}$ be the sequence of maps that define the zooms near the parameter $s_{\ell}$ for the measure $\mu^{r+1}_{\lambda_{1}}$ corresponding to the transversely prerepelling critical point at $x_{1}$, as described before Lemma 4.1. From [Ga2, Proposition 3.2] we know that there is a measure $\mu\neq 0$ (which can be described explicitly) such that
(4.11) $d^{nr}\,(r_{n})^{*}(\mu^{r+1}_{\lambda_{1}})\to\mu,$
uniformly as $n\to\infty$. On the other hand, it follows from (4.10) that the functions $U_{n}\circ r_{n}$ also converge uniformly to 0. We conclude that $d^{n}\,(r_{n})^{*}T_{c_{r+1}}\to 0$ weakly, and therefore
@@
Therefore, $\mu^{1}_{\lambda_{1}}$ cannot be proportional to $\mu^{r+1}_{\lambda_{1}}$, contradicting (4.8). This completes the proof of Theorem 1.8.
@@
The proof would simplify if we knew that the $(r+1)$-th critical point $c_{r+1}$ were preperiodic to a parabolic cycle, instead of only in the basin of that cycle. We could employ ideas that appeared in [GNY, Proposition 5.2] (written for measures in the dynamical plane) or the proof of [FG3, Proposition 5.12] (for $1$-parameter families of polynomials and marked points). But the critical point $c_{r+1}$ is sometimes the only “free” critical point and must therefore have an infinite orbit in the presence of a parabolic cycle.
@@
We then repeat the argument with the families of curves over the positive-dimensional components in the complement of $V\setminus U$; we work with irreducible components so that the generic fibers of these families will be (geometrically) irreducible. Note that each family we treat is defined over $\overline{\mathbb{Q}}$. We increase the bound $N$ and decrease $\epsilon$ as needed. We are left with a finite set $S$ in $V(\overline{\mathbb{Q}})$ containing all of the irreducible PCF-special curves with degree $\leq D$ as well as all curves in $V$ with infinitely many points of critical height $\frac{\varepsilon}{\ell+1}\max\{1,h_{\mathrm{Ch}}(\lambda)\}$ (or $\frac{\varepsilon}{\ell+1}$ over complex parameters). This proves Theorems 1.2 and 1.6 for the curves in $V$. If for any of the curves $C_{\lambda}$ in this leftover finite set $S\subset V(\overline{\mathbb{Q}})$, we have $N<|\Sigma_{\lambda,\varepsilon}|<\infty$, we just increase the bound $N$ as needed so that Theorem 1.5 and 1.2 hold for all curves in $V$. We repeat the argument for all irreducible components $V$ of $\mathrm{Ch}(\mathrm{M}_{d},1,D)$ with geometrically irreducible generic fiber and adjust the constants $\varepsilon,~N$ as needed. This proves Theorems 1.1, 1.2, 1.5 and 1.6 for curves.
Now fix $r>1$ and suppose we know the conclusion of Theorems 1.1, 1.2, 1.5, and 1.6 for subvarieties in $\mathrm{M}_{d}$ of dimension <r<r. Consider any family $\mathcal{X}\to V$ of $r$-dimensional varieties in $\mathrm{M}_{d}$ over an irreducible and quasiprojective $V\subset\mathrm{Ch}(\mathrm{M}_{d},r,D)$ defined over $\overline{\mathbb{Q}}$ and of dimension $\ell\geq 1$ for which the generic fiber of $\mathcal{X}\to V$ is geometrically irreducible.
@@
for $\lambda\in U(\mathbb{C})\setminus U(\overline{\mathbb{Q}})$, lie in a proper algebraic subvariety $Z_{(\lambda)}\subset X_{\lambda}$ which has $\leq N_{1}$ irreducible components, each with degree $\leq B_{1}$, for all $\lambda\in U(\mathbb{C})$. Note that $\Sigma^{r}_{\lambda,\epsilon}$ is not necessarily Zariski dense in $Z_{(\lambda)}$, so we apply the induction hypotheses (specifically, the constants from each of the statements of Theorems 1.1, 1.2, 1.5, 1.6 for subvarieties of dimension <r<r) to each irreducible component $Y_{(\lambda)}$ of $Z_{(\lambda)}$, for all $\lambda\in U(\mathbb{C})$. Thus, by reducing $\varepsilon$ if needed, we conclude that the Zariski-closure of $\Sigma^{r}_{\lambda,\varepsilon}\cap Y_{(\lambda)}$ has uniformly bounded number of components, each with uniformly bounded degree for all $\lambda\in U(\mathbb{C})$ . Further, adjusting these uniform bounds if needed, using the induction hypothesis, we can arrange that they also bound the geometry of the Zariski closure of $\Sigma^{r}_{\lambda,\varepsilon^{\prime}}\cap Y_{(\lambda)}\subset\Sigma^{r}_{\lambda,\varepsilon}\cap Y_{(\lambda)}$ for all $0\leq\varepsilon^{\prime}<\varepsilon$ (and in particular of $Y_{(\lambda)}\cap\mathrm{PCF}_{d}$). Combined with the initial bound of $N_{1}$ on the number of components of $Z_{(\lambda)}$, repeating this procedure provides uniform bounds on the Zariski closures of $\Sigma^{r}_{\lambda,\varepsilon^{\prime}}$ for all $\varepsilon^{\prime}<\varepsilon$ and of $\overline{X_{\lambda}\cap\mathrm{PCF}_{d}}$ for all $\lambda\in U(\mathbb{C})$.